# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 11 Offline Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.139139139In causal inference, “outcome” refers specifically to the response variable $Y$ (or the potential outcome $Y(d)$ in the Rubin framework). The RL/MDP framework has no single corresponding concept called “outcome”; the relevant quantities are the reward $R_{t}$, the next state $S_{t+1}$, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k}$, or the value function $V^{\pi}(s)$. When economists refer to the “outcome model” $\mathbb{E}[Y\mid X,D]$, the RL analogue is the action-value function $Q(s,a)$. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@
Unobserved demand shocks that affect both a retailer’s pricing algorithm and consumer demand create endogeneity: the observed correlation between price and demand conflates the causal effect with the confounding effect. An RL agent trained on such observational data converges to a biased policy that systematically overestimates the revenue from high prices.
This chapter formalizes the confounded MDP, presents backdoor-adjusted off-policy evaluation as the key algorithmic solution, and demonstrates its practical consequences through a simulation study in a confounded chain MDP. For comprehensive surveys of the rapidly growing causal RL literature, including causal representation learning, counterfactual policy optimization, transfer, and fairness, I refer readers to Deng et al. (2023) and da Costa Cunha et al. (2025).
@@ 11.1 The Confounded MDP
Standard MDPs assume that the state $s_{t}$ is fully observed and that the agent’s policy depends only on $s_{t}$. In the online algorithms of earlier chapters, these assumptions held by construction: the agent executed its own policy in a known environment, so no unobserved variable could simultaneously influence action selection and transitions. When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.140140140The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
A confounded MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{U},P,r,\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite action space, $\mathcal{U}$ is a space of unobserved confounders, $\gamma\in[0,1)$ is a discount factor, and the dynamics are governed by structural equations:
@@
The behavioral (logging) policy $\mu$ depends on the unobserved confounder $U_{t}$ through Equation (80). An evaluation policy $\pi(a\mid s)$ depends only on the observed state.141141141A confounded MDP differs from a partially observed MDP (POMDP). In a POMDP, the agent knows its observation is incomplete and optimizes accordingly; the hidden state affects what the agent sees but the analyst models this explicitly. In a confounded MDP, the hidden variable $U_{t}$ distorts the behavioral policy $\mu$ without the analyst’s knowledge or control, so the core challenge is not planning under partial observability but rather identifying interventional quantities from data generated by a policy that depended on variables the analyst cannot observe.
The separation between the behavioral policy $\mu$ and the evaluation policy $\pi$ is the formal signature of offline RL. In the online algorithms of Sections 35, the behavior and target policies were identical (on-policy methods) or related by a known exploration mechanism ($\epsilon$-greedy, softmax); here, $\mu$ is an unknown function of unobserved variables.
The key distinction is between observational and interventional transition dynamics. Under the behavioral policy $\mu$, which depends on $U_{t}$, the conditional distribution $P(s^{\prime}\mid s,a)$ obtained by conditioning on the observed action does not equal the interventional distribution obtained by setting the action externally:
@@
The left-hand side conditions on the event $\{A_{t}=a\}$, which carries information about $U_{t}$ through the behavioral policy $\mu$. The right-hand side intervenes on $A_{t}$, severing the dependence between action and confounder. When $U_{t}$ also affects $S_{t+1}$ through Equation (82), the conditioning introduces a spurious association between $a$ and $s^{\prime}$.
@@
When an analyst ignores confounding and uses the observational transition probabilities $P(s^{\prime}\mid s,a)$ in place of $P(s^{\prime}\mid s,\operatorname{do}(a))$, the resulting value estimates are biased.
@@
Let $\hat{V}^{\pi}_{\text{naive}}$ denote the value function obtained by solving the Bellman equation with observational transitions $P(s^{\prime}\mid s,a)$, and let $V^{\pi}$ denote the true value function under interventional transitions $P(s^{\prime}\mid s,\operatorname{do}(a))$. In a confounded MDP where $P(s^{\prime}\mid s,a)\neq P(s^{\prime}\mid s,\operatorname{do}(a))$ for some $(s,a,s^{\prime})$, the naive estimator is biased:
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.
The bias arises because the importance-sampling ratio $\pi(a\mid s)/\mu(a\mid s)$ does not correctly reweight the data. The true behavioral policy is $\mu(a\mid s,u)$, but since $u$ is unobserved, the analyst can only estimate the marginal propensity $\mu(a\mid s)=\mathbb{E}_{U}[\mu(a\mid s,U)\mid s]$. Reweighting by the marginal propensity does not eliminate the confounding bias.142142142This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.143143143A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@
Substituting Equation (86) into the causal Bellman operator (Equation 84) yields an identified, consistent estimator of $V^{\pi}$.144144144While Theorem 6 assumes $\mathbf{Z}_{t}$ blocks all backdoor paths, identification is also possible when $\mathbf{Z}_{t}$ are merely proxies (effects) of the unobserved confounder. This “proximal” approach to RL employs the theory of negative controls to recover interventional values without satisfying the backdoor criterion (Bennett and Kallus, 2021).
@@ 11.2 Backdoor-Adjusted Off-Policy Evaluation
Off-policy evaluation under confounding is precisely an average treatment effect (ATE) estimation problem in a dynamic setting (Bannon et al., 2020). The behavioral policy $\mu$ corresponds to the treatment assignment mechanism, the target policy $\pi$ to the counterfactual treatment regime, and the policy value $V^{\pi}$ to the expected potential outcome $\mathbb{E}[Y(d)]$ under the Rubin framework. Importance sampling in OPE corresponds to inverse probability weighting (IPW); doubly robust OPE corresponds to the augmented IPW (AIPW) estimator of Robins et al. (1994). The key challenge, distribution shift between $\mu$ and $\pi$, is the analogue of the overlap (positivity) condition in causal inference.
In the online setting of earlier chapters, distribution shift did not arise: the data distribution tracked the current policy throughout training, so the state-action frequencies under which value estimates were formed matched the frequencies under which the policy would be deployed.
Theorem 6 yields a concrete estimation procedure. Given logged data $\{(s_{t},a_{t},z_{t},r_{t},s_{t+1})\}_{t=1}^{N}$ collected under behavioral policy $\mu$, where $z_{t}$ is an observed proxy for the confounder:

### 15 Mar 2026 v2

@@ 13 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.200200200See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@
Unobserved demand shocks that affect both a retailer’s pricing algorithm and consumer demand create endogeneity. The observed correlation between price and demand conflates the causal effect with the confounding effect. An RL agent trained on such observational data converges to a biased policy that systematically overestimates the revenue from high prices.
This chapter formalizes the confounded MDP, develops four identification strategies for recovering interventional quantities from observational data (backdoor adjustment, front-door adjustment, instrumental variables, and proximal causal inference), and demonstrates their practical consequences through a unified simulation study. For comprehensive surveys of the rapidly growing causal RL literature, including causal representation learning, counterfactual policy optimization, transfer, and fairness, I refer readers to Deng et al. (2023) and da Costa Cunha et al. (2025).
13.1 From Partial Observability to Causal Structure
Before formalizing confounded MDPs, it is useful to distinguish partial observability from confounding, since both involve hidden variables but create fundamentally different challenges. A partially observable MDP (POMDP) augments the standard MDP with an observation function. The POMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,O,r,\gamma)$, where $\mathcal{O}$ is a finite observation space and $O$ is the observation function.
$O(o\mid s^{\prime},a)=P(O_{t}=o\mid S_{t}=s^{\prime},A_{t-1}=a).$ (101)
The agent does not observe $s_{t}$ directly but instead receives $o_{t}\sim O(\cdot\mid s_{t},a_{t-1})$ and maintains a belief state $b_{t}\in\Delta(\mathcal{S})$
$b_{t}(s)=P(S_{t}=s\mid o_{1},a_{1},\ldots,o_{t}),$ (102)
which is updated via Bayesian filtering at each step. The belief MDP, whose state space is $\Delta(\mathcal{S})$, is itself a fully observable (continuous-state) MDP, so standard value iteration applies in principle, though computation is intractable in general.
da Costa Cunha et al. (2025) organize sequential decision problems with hidden variables into a hierarchy: standard MDP (full observability, no confounding), POMDP (partial observability, no confounding), confounded MDP (full observability, confounding), and causal POMDP (both). The key distinction is epistemic versus identificational. In a POMDP, the hidden state is a modeling challenge: the agent acknowledges incomplete information and plans accordingly via the belief state, analogous to Kalman or Hamilton filtering in econometrics. In a confounded MDP, the hidden variable is an identification challenge: standard estimators silently produce biased results, analogous to endogeneity and omitted variable bias.
@@ 13.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.201201201The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
A confounded MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{U},P,r,\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite action space, $\mathcal{U}$ is a space of unobserved confounders, $\gamma\in[0,1)$ is a discount factor, and the dynamics are governed by structural equations
@@
The behavioral (logging) policy $\mu$ depends on the unobserved confounder $U_{t}$ through Equation (104). An evaluation policy $\pi(a\mid s)$ depends only on the observed state.
Unlike the online algorithms of Sections 47, where behavior and target policies were identical or related by a known exploration mechanism, here $\mu$ is an unknown function of unobserved variables.
Because $\mu$ depends on $U_{t}$, conditioning on $\{A_{t}=a\}$ carries information about the confounder, so the observational and interventional transitions diverge:
@@
Let $\hat{V}^{\pi}_{\text{naive}}$ denote the value function obtained by solving the Bellman equation with observational transitions $P(s^{\prime}\mid s,a)$, and let $V^{\pi}$ denote the true value function under interventional transitions $P(s^{\prime}\mid s,\operatorname{do}(a))$. In a confounded MDP where $P(s^{\prime}\mid s,a)\neq P(s^{\prime}\mid s,\operatorname{do}(a))$ for some $(s,a,s^{\prime})$, the naive estimator is biased.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.202202202This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.203203203A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@
Substituting Equation (110) into the causal Bellman operator (Equation 108) yields an identified, consistent estimator of $V^{\pi}$.
@@ 13.3 Backdoor-Adjusted Off-Policy Evaluation
Off-policy evaluation under confounding is an average treatment effect estimation problem in a dynamic setting (Bannon et al., 2020): $\mu$ is the treatment assignment mechanism, $\pi$ the counterfactual regime, importance sampling corresponds to inverse probability weighting, and doubly robust OPE corresponds to the AIPW estimator of Robins et al. (1994).
Theorem 8 yields a concrete estimation procedure. Given logged data $\{(s_{t},a_{t},z_{t},r_{t},s_{t+1})\}_{t=1}^{N}$ collected under behavioral policy $\mu$, where $z_{t}$ is an observed proxy for the confounder.
@@
13.4 Alternative Identification Strategies
When no backdoor variable is available, three alternative identification strategies apply, each with a direct econometric analogue. Figure 25 displays the causal graph for each.
Figure 25: Causal graphs for three identification strategies in confounded MDPs. Gray dashed nodes are unobserved; dashed edges involve unobserved variables. (a) Front-door criterion with mediator $M_{t}$. (b) Instrumental variables with exogenous instrument $Z_{t}$. (c) Proximal causal inference with proxies $W_{t}^{(1)},W_{t}^{(2)}$.
13.4.1 Front-Door Criterion
The front-door criterion applies when $A_{t}$ affects $S_{t+1}$ only through an observed mediator $M_{t}$. Three conditions are required: (i) $M_{t}$ intercepts all directed paths from $A_{t}$ to $S_{t+1}$, (ii) no unblocked backdoor path exists from $A_{t}$ to $M_{t}$, and (iii) $A_{t}$ blocks all backdoor paths from $M_{t}$ to $S_{t+1}$. When satisfied (Pearl, 2009):
$P(s^{\prime}\mid s,\operatorname{do}(a))=\sum_{m}P(m\mid a)\sum_{a^{\prime}}P(s^{\prime}\mid s,m,a^{\prime})\,P(a^{\prime}\mid s).$ (112)
The first factor is unconfounded; the inner sum adjusts for confounding on the $M_{t}\to S_{t+1}$ link by averaging over the observational action distribution. This is the sequential analogue of mediation analysis (da Costa Cunha et al., 2025).
The causal RL survey of da Costa Cunha et al. (2025) illustrates the front-door criterion with a mobile wellness intervention. A health app (action $A_{t}$) aims to reduce patient cortisol levels (outcome $S_{t+1}$), but unobserved health consciousness $U_{t}$ confounds the relationship because health-conscious individuals are both more likely to adopt the app and more likely to have low cortisol regardless. The app affects cortisol only through an observed mediator, supplement adherence $M_{t}$. Applying Equation (112) recovers the causal effect without observing health consciousness.
13.4.2 Instrumental Variables
When neither backdoor nor front-door variables are available, instrumental variable methods can identify causal effects. An instrument $Z_{t}$ must satisfy two conditions: it affects the action $A_{t}$ (relevance) and its effect on $S_{t+1}$ is channeled entirely through $A_{t}$ (exclusion restriction). Liao et al. (2024) formalize this as a Confounded MDP with Instrumental Variables (CMDP-IV), where transitions take the form $S_{t+1}=F^{*}(S_{t},A_{t})+\epsilon_{t}$ with unobserved confounders $\epsilon_{t}$ affecting both the behavior policy and transitions. The transition function $F^{*}$ is recovered from a conditional moment restriction:
$\mathbb{E}[S_{t+1}-F^{*}(S_{t},A_{t})\mid Z_{t},S_{t}]=0.$ (113)
For a binary action and binary instrument, the Wald estimator provides a closed-form solution. Let $\beta$ denote the causal effect of promoting (action $a=0$) on the transition probability.
$\beta=\frac{P(s^{\prime}\mid s,Z{=}1)-P(s^{\prime}\mid s,Z{=}0)}{P(A{=}0\mid s,Z{=}1)-P(A{=}0\mid s,Z{=}0)},$ (114)
where the numerator is the reduced-form effect of the instrument on the transition and the denominator is the first-stage effect on treatment uptake. The interventional transition is then recovered from any instrument value $z$ via $P(s^{\prime}\mid s,\operatorname{do}(a{=}0))=P(s^{\prime}\mid s,Z{=}z)+\beta\cdot(1-P(A{=}0\mid s,Z{=}z))$. Their IV-aided Value Iteration algorithm applies this moment restriction at each state to estimate $F^{*}$, then runs standard value iteration on the estimated model.
Liao et al. (2024) illustrate with neonatal intensive care unit (NICU) assignment. A hospital must decide whether to admit each newborn to the NICU (action $A_{t}$), and this decision is confounded by unobserved severity indicators that affect both the admission decision and the infant’s health trajectory. Differential travel time from the hospital to a specialty care provider serves as an instrument $Z_{t}$. Relevance holds because longer travel times discourage NICU referrals, shifting the admission probability. The exclusion restriction holds because travel time affects infant health outcomes only through the admission decision, not directly. The conditional moment restriction (Equation 113) uses variation in travel time across hospitals to trace out the causal effect of NICU admission on health transitions, recovering the transition function $F^{*}$ that standard regression conflates with the unobserved confounder.
13.4.3 Proximal Causal Inference
When confounders are truly latent but proxy variables, noisy correlates of the confounder, are available, proximal causal inference provides identification. Bennett and Kallus (2021) adapt this to sequential settings. The analyst observes two proxies: a treatment-side proxy $W_{t}^{(1)}$ and an outcome-side proxy $W_{t}^{(2)}$, both conditionally independent given the latent confounder $U_{t}$. Identification proceeds through a bridge function $h$ that solves a conditional moment equation linking the two proxies to the interventional quantity:
$\mathbb{E}[V(S_{t+1})\mid W_{t}^{(1)},S_{t},A_{t}]=\mathbb{E}[h(W_{t}^{(2)},S_{t},A_{t})\mid W_{t}^{(1)},S_{t},A_{t}].$ (115)
The bridge function $h$ is estimated by solving this integral equation (which reduces to a linear system in the discrete case), and the causal effect is recovered by marginalizing over the outcome proxy distribution.
$P(s^{\prime}\mid s,\operatorname{do}(a))=\sum_{w_{2}}h(w_{2},s,a)\,P(W^{(2)}{=}w_{2}\mid s).$ (116)
Two bridge functions, analogous to inverse propensity scores and Q-functions, yield a doubly robust estimator of the policy value that is $\sqrt{n}$-consistent without ever observing $U_{t}$. Bennett and Kallus (2021) demonstrate with a sepsis management simulator in which physicians choose among fluids, vasopressors, and antibiotics at each decision point. The patient’s diabetes status is the latent confounder, partially censored in the medical record so that 20% of diabetic patients appear as non-diabetic in the data. Because the physician observes the true diabetes status when prescribing but the analyst’s dataset contains the censored version, the behavioral policy depends on a variable the analyst cannot fully recover. Previous clinical observations $O_{t-1}$ (prior lab values and vitals) serve as the treatment-side proxy $W_{t}^{(1)}$ because they correlate with diabetes status and influence the physician’s current prescribing. Current clinical observations $O_{t}$ serve as the outcome-side proxy $W_{t}^{(2)}$ because they reflect diabetes status and predict future health transitions. In their experiments, the proximal estimator correctly identified which evaluation policy improved over the behavioral policy in 82–100% of test cases, while naive estimators that ignored confounding and standard MDP estimators that assumed full observability both achieved 0% accuracy.